Express v=(1,-2,5) in R3 as a linear combination of the vectors U1=(1,1,1), U2=(1,2,3), U3=(2,-1,1)


1
Expert's answer
2022-01-10T15:10:41-0500

Let "v=xu_1+yu_2+zu_3." Then

"(1,-2,5)=x(1,1,1)+y(1,2,3)+z(2,-1,1)=(x+y+2z,x+2y-z,x+3y+z),"

and hence we get the following system


"\begin{cases}
x+y+2z=1\\
x+2y-z=-2\\
x+3y+z=5
\end{cases}"


which is equivalent (after substracting from the second row the first row, from the third row the second row) to


"\begin{cases}
x+y+2z=1\\
y-3z=-3\\
y+2z=7
\end{cases}"


and (after substracting from the third row the second row) to


"\begin{cases}
x+y+2z=1\\
y-3z=-3\\
5z=10
\end{cases}"


It follows that "z=2,\ y =3z-3=6-3=3," and "x=1-y-2z=1-3-4=-6."


We conclude that "v=-6u_1+3u_2+2u_3."